Question 12

The production function expressed in the table below represents the number of tacos that can be made at a taco truck in a given day based on the number of workers. Calculate the marginal product of adding the 3rd worker.

# of Workers 1 2 3 4 5
# of Tacos Produced 105 220 350 465 515
Answer #1

Marginal product of the labor = Change in the output / change in the number of labor.

350 - 220 / 3-2 = 130 / 1 = 130.

the marginal product of the third labor in the firm is 130.